AS2650(1L) AS2650(1L) NAME as2650−crossassemblerformicrocomputers SYNOPSIS as2650[-pcpu][-llistfile][-ohexfile][-d][-ssymbolfile]input DESCRIPTION Theas2650commandassemblestheinputfileintoatextoutputfilerepresentingtheprogram memoryforamicrocomputer. Options -pcpuIgnoredbythisprogram. -llistfile Outputafileformatedwiththememoryaddress,data,andsourceinputlines. -ohexfile Outputthememorydataimageinaformacceptedbymostpromprogrammers. -hhexfile Sameas−o. −dSavetheintermediatedatafile(seeFILE)andaborttheexecutionattheendofprocessing. -ssymbolfile Printthesymboltablevaluesandnames,oneperlineinthespecifiedfile. FILES /usr/tmp/frtXXXXXX SEEALSO FrankensteinCrossAssemblersUsersManual(filebase.doc)Appendixforas2650Frankenstein Assembler(fileas2650.doc) NOTES Thereisonlyoneinputfile. 1